OpenStudy (anonymous):

Simplify by factoring. Assume that all expressions under radicals represent nonnegative numbers. 3 sqrt 243 x^8

OpenStudy (saifoo.khan):

\[\sqrt[3]{243x^8}\]Like that?

OpenStudy (anonymous):

27x^4 sqrt 3

OpenStudy (anonymous):

3 * sqrt ( 243 * x^8) = 3 * sqrt ( 9 * 9 * 3 * x^4 * x^4) = 3 * 9 * x^4 * sqrt (3)

OpenStudy (anonymous):

thus 27x^4 sqrt 3

OpenStudy (anonymous):

Thank you all!